# Liam Welsh

### v2

@@ 3 OC Market Structure and RL Formulation
In this section we detail the OC market model setting, and resulting game, that we work in. The model setting is a discrete-time and finite-agent extension of the model originally presented in continuous time (with one or two agents) in Welsh and Jaimungal, (2024). Throughout, we work on a completed and filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T_{L}]},\mathbb{P})$, where the OC price is denoted $S=\left(S_{t}\right)_{t\in[0,T_{L}]}$, and $(\mathcal{F}_{t})_{t\in[0,T_{L}]}$ is the natural filtration generated by $S$. We assume the market is populated with $N$-many agents, where agents are labeled with $i\in{\mathfrak{N}}$ and that there are $L$ many compliance periods, with compliance dates (i.e the date on which firms must submit OCs or pay penalties for the prior period) occur on the dates in the set $\{T_{l}:l\in{\mathfrak{L}}\}$, where ${\mathfrak{L}}:=\{1,\dots,L\}$. In Section 2.2 time was discretized such that time evolved as $t\rightarrow t+1$. In the sequel, we use a time grid such that all compliance periods are discretized into equally spaced partitions and define ${\mathfrak{T}}:=\{t_{0},t_{1},\dots,t_{K}=T_{L}\}$ where $\Delta t=t_{k+1}-t_{k}$. OCs acquired prior to compliance date $T_{l}$ may be used for regulatory purposes on $T_{l}$. On a compliance date $T_{l}$, the regulator imposes a cost of
@@
where $Z_{t_{k}}\sim N(0,1)$ and $U_{t_{k},n}\sim\mathcal{U}(0,1)$, for all $t_{k}\in{\mathfrak{T}}$, $t_{k}\in[T_{l-1},T_{l})$ and $l\in{\mathfrak{L}}$. This evolution is a discretely sampled Brownian bridge between compliance dates, where the bridge is pinned to $\mathfrak{p}$} at the end of each compliance period. This pinning prevents potential arbitrage opportunities so that OCs cannot trader at a discount or premium with respect to the penalty near a compliance date.
@@ 8   Perform soft update of target network parameters $\tilde{\alpha}$ using(19);
$\sum_{l\in{\mathfrak{L}}}G_{l,i}(X_{T_{l},i})$ $=L\,{\mathfrak{p}}({\mathfrak{R}}-X_{0,i})_{+}+\sum_{t=0}^{T_{L}-1}F_{t_{k}}\;g(X_{t_{k+1},i},X_{t_{k},i})$
@@ 4.1 Four Agent Setting
Figure 3 shows the agents’ inventories. The shaded region illustrates the quantile bands and the solid lines the average inventories. As observed, no agent fulfills the full requirement of 25 OCs in either period; consequently, all agents incur a partial penalty.
@@
Figure 4 displays the agent trade rates (left panel) and generation probabilities (right panel). All agents take part in generating OCs to varying degrees, with Agents 1 (blue), 3 (green), and 4 (red) all generating substantially more than Agent 2 (orange). Agents 2, 3, and 4 typically trade at positive rates, while Agent 1, the largest agent, takes the role of the market clearing agent, selling a partial amount of their inventory which also partially covers their cost of generating.
@@ 5 Conclusion
In this work, we developed a finite-agent GHG OC model based on the Canadian market and approximated the corresponding Nash equilibrium using the deep RL methodology Nash-DQN. We applied our methodology to a four player and an eight player settings, and in doing so we demonstrated that regulated firms who actively participate in the market (e.g. either through trading and/or generating OCs) achieve non-trivial financial savings when abiding by the Nash equilibrium strategy. Performance is measured with respect to agents’ mean PnLs and TEs against their benchmarks (i.e. full penalty value). As more agents were added to the market model, we found that a greater percentage of OCs were generated, with respect to the total amount required to offset every agents’ penalty. While computing a Nash equilibrium is an NP-hard problem, we are able to efficiently approximate it in our finite-agent settings. Further computational savings are achieved by allowing agents to share DNNs when generation capabilities and OC requirements are identical. In our methodology we also include a soft market clearing condition through an additional term on the loss function, thus creating a more realistic market environment.
@@
The methodology and analysis here can aid regulators to understand how regulated firms may behave in these markets and can lead to legislative improvements in the market. Given the drastically increasing and recurring consequences of man-made climate change, the importance for firms to participate in GHG OC markets and either reduce GHG emissions or implement GHG capturing projects is vital to reduce these consequences. By employing modern statistical learning techniques to solve for the optimal policies we are able to demonstrate the financial viability of active participation in GHG OC markets for firms, as well as the societal benefits these markets have in mitigating the impacts of climate change.
@@
Both climate finance and RL (and more generally machine learning) are flourishing areas of research, hence there are many open problems that intersect the two. Within the current framework, there remain open problems that are worthwhile investigating. First, this paper’s goal was to illustrate the viability of deploying Nash-DQN in this offset credit marketing setting, and we did not calibrate our model to real data. This is because the updated Canadian market is still verifying submitted projects and no OCs have been generated, and further calibrating model parameters in a highly non-trivial problem that requires (often proprietary) knowledge of a firm’s internal finances. Future work can bridge this gap by calibrating and tuning our model to real data, once it is available. Another potential avenue to take is to create a principal agent game version of our market model, such that the principal agent takes the place of a market regulator who has their own goals which the other agents (e.g. firms) must adapt to. Campbell et al., (2021) and Ivanov et al., (2024) have demonstrated this is a viable area of study, developing a scalable framework using contract theory and RL methods. Another worthwhile pursuit is to introduce stochastic OC requirements into the model. Currently, we assume the agents’ requirements are deterministic and exogenous, which may not be true in the real world. Extending the methodology to include stochastic requirements would provide further complexity and realism to the model. Finally, while our model includes price impact, the OC price is given exogenously. Endogenizing the OC price within the current setting would be a very interesting avenue to explore. Given the lack of extant literature for price endogenization in model-free RL, this has the potential to be adapted into many other types of models.

### v3

@@ 3 OC Market Structure and RL Formulation
In this section we detail the OC market model setting, and resulting game, that we work in. The model setting is a discrete-time and finite-agent extension of the model originally presented in continuous time (with one or two agents) in Welsh and Jaimungal (2024). Throughout, we work on a completed and filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T_{L}]},\mathbb{P})$, where the OC price is denoted $S=\left(S_{t}\right)_{t\in[0,T_{L}]}$, and $(\mathcal{F}_{t})_{t\in[0,T_{L}]}$ is the natural filtration generated by $S$. We assume the market is populated with $N$-many agents, where agents are labeled with $i\in{\mathfrak{N}}$ and that there are $L$ many compliance periods, with compliance dates (i.e the date on which firms must submit OCs or pay penalties for the prior period) occur on the dates in the set $\{T_{l}:l\in{\mathfrak{L}}\}$, where ${\mathfrak{L}}:=\{1,\dots,L\}$. In Section 2.2 time was discretized such that time evolved as $t\rightarrow t+1$. In the sequel, we use a time grid such that all compliance periods are discretized into equally spaced partitions and define ${\mathfrak{T}}:=\{t_{0},t_{1},\dots,t_{K}=T_{L}\}$ where $\Delta t=t_{k+1}-t_{k}$. OCs acquired prior to compliance date $T_{l}$ may be used for regulatory purposes on $T_{l}$. On a compliance date $T_{l}$, the regulator imposes a cost of
@@
where $Z_{t_{k}}\sim N(0,1)$ and $U_{t_{k},n}\sim\mathcal{U}(0,1)$, for all $t_{k}\in{\mathfrak{T}}$, $t_{k}\in[T_{l-1},T_{l})$ and $l\in{\mathfrak{L}}$. This evolution is a discretely sampled Brownian bridge between compliance dates, where the bridge is pinned to $\mathfrak{p}$} at the end of each compliance period. This pinning prevents potential arbitrage opportunities so that OCs cannot be traded at a discount or premium with respect to the penalty near or on a compliance date.
@@ 8   Perform soft update of target network parameters $\tilde{\alpha}$ using(19);
$\sum_{l\in{\mathfrak{L}}}G_{l,i}(X_{T_{l},i})$ $=L\,{\mathfrak{p}}({\mathfrak{R}}-X_{0,i})_{+}+\sum_{t=0}^{T_{L}-\Delta t}F_{t_{k}}\;g(X_{t_{k+1},i},X_{t_{k},i})$
@@ 4.1 Four Agent Setting
Figure 3 shows the agents’ inventories. The shaded region illustrates the quantile bands and the solid lines the average inventories. As observed, no agent fulfills the full requirement of 25 OCs in either period; consequently, all agents incur a partial penalty. The sudden drop in inventory all agents experience on compliance dates is caused by agents submitting their OCs to the regulator to reduce/eliminate their compliance penalty. If an agent had an inventory exceeding their requirement at a compliance date, then they are able to carry over their excess inventory into the next period.
@@
Figure 4 displays the agent trade rates (left panel) and generation probabilities (right panel). All agents take part in generating OCs to varying degrees, with Agents 1 (blue) and 4 (red) all generating substantially more than Agent 2 (orange) and Agent 3 (green). Agent 1 generates with a probability of approximately 50% throughout both periods, experiencing a minor increase in generation probability in the second period compared to the first period. This can be explained as the agent having more urgency to generate OCs as they enter the terminal compliance period. Agents 2, 3, and 4 typically trade at positive rates, while Agent 1, the largest agent, takes the role of the market clearing agent, selling a partial amount of their inventory which also partially covers their cost of generating.
@@ 5 Conclusion
In this work, we developed a finite-agent GHG OC model based on the Canadian market and approximated the corresponding Nash equilibrium using the deep RL methodology Nash-DQN. We applied our methodology to a four player and an eight player settings, and in doing so we demonstrated that regulated firms who actively participate in the market (e.g. either through trading and/or generating OCs) achieve non-trivial financial savings when abiding by the Nash equilibrium strategy. Performance is measured with respect to agents’ mean PnLs and TEs against their benchmarks (i.e. full penalty value). As more agents were added to the market model, we found that a greater percentage of OCs were generated, with respect to the total amount required to offset every agents’ penalty. These experiments demonstrate to regulators that the current market setting provides a fertile environment for firms to invest in OC generation, thus capturing atmospheric $\text{CO}_{2}\text{e}$ and reducing overall emissions. While some firms actively generate OCs and others do not, this represents the broad spectrum of firm behaviour in the market, with some companies placing higher emphasis on environmental governance than others. Regulators may wish to create a less hospitable market environment for non-OC generating firms, and may experiment with design using the Nash-DQN framework we present. While computing a Nash equilibrium is an NP-hard problem, we are able to efficiently approximate it in our finite-agent settings. Further computational savings are achieved by allowing agents to share DNNs when generation capabilities and OC requirements are identical. In our methodology we also include a soft market clearing condition through an additional term on the loss function, thus creating a more realistic market environment.
@@
The methodology and analysis here can aid regulators to understand how regulated firms may behave in these markets and can lead to legislative improvements in the market by providing a computational and experimental framework in which new policies and market designs can be effectively tested. The adaptability of the Nash-DQN framework will allow regulators to efficiently test various market perturbations. By analyzing firm behaviour across many experimental settings, regulators can update the market in ways to increase OC generation, OC liquidity, and/or market participation. Given the drastically increasing and recurring consequences of man-made climate change, the importance for firms to participate in GHG OC markets and either reduce GHG emissions or implement GHG capturing projects is vital to reduce these consequences. By employing modern statistical learning techniques to solve for the optimal policies we are able to demonstrate the financial viability of active participation in GHG OC markets for firms, as well as the societal benefits these markets have in mitigating the impacts of climate change.
@@
Both climate finance and RL (and more generally machine learning) are flourishing areas of research, hence there are many open problems that intersect the two. Within the current framework, there remain open problems that are worthwhile investigating. First, this paper’s goal was to illustrate the viability of deploying Nash-DQN in this offset credit marketing setting, and we did not calibrate our model to real data. This is because the updated Canadian market is still verifying submitted projects and no OCs have been generated, and further calibrating model parameters in a highly non-trivial problem that requires (often proprietary) knowledge of a firm’s internal finances. Future work can bridge this gap by calibrating and tuning our model to real data, once it is available. Another potential avenue to take is to create a principal agent game version of our market model, such that the principal agent takes the place of a market regulator who has their own goals which the other agents (e.g. firms) must adapt to. Campbell et al. (2021) and Ivanov et al. (2024) have demonstrated this is a viable area of study, developing a scalable framework using contract theory and RL methods. Another worthwhile pursuit is to introduce stochastic OC requirements into the model. Currently, we assume the agents’ requirements are deterministic and exogenous, which may not be true in the real world. Extending the methodology to include stochastic requirements would provide further complexity and realism to the model. Finally, while our model includes price impact, the OC price is given exogenously. Endogenizing the OC price within the current setting would be a very interesting avenue to explore. Given the lack of extant literature for price endogenization in model-free RL, this has the potential to be adapted into many other types of models.